OpenStudy (deercult):

53. The measure of the supplement of an angle is 20 degrees more than three times the measure of the original angle. Find the measures of the angles. So far I have: s = 3a + 20 s + a = 180

OpenStudy (deercult):

I always get lost in the middle of working these out ;~;

OpenStudy (joshtray):

Have you learnt simultaneous equations?

OpenStudy (joshtray):

Because that is what you will need for this question

OpenStudy (deercult):

Yeah! I'm practicing them now. I just keep getting lost on where to go after writing an equation down.

OpenStudy (deercult):

RIP.

OpenStudy (deercult):

AGH I feel dumb, I swear every time I post these I have a lightbulb moment two minutes later

TheSmartOne (thesmartone):

we can check your work then just to make sure you're on the right path :3

OpenStudy (deercult):

Wait, no, that's not gonna help me. I was doing it wrong and got s-20/3 = a, not much help. Should I make that into the 180 equation? Like, s + (s-20/3) = 180, or am I just digging myself into a deeper hole there?

OpenStudy (joshtray):

Well, that could work

OpenStudy (joshtray):

Carry on

TheSmartOne (thesmartone):

you have the two equations correct s = 3a + 20 s + a = 180 You can solve them by substitution

OpenStudy (joshtray):

That is what he has started with @TheSmartOne

OpenStudy (deercult):

Yea! I reversed the first equation, sorta, so I had one variable to work with. I got an s answer so far of 330, so I'm about to plug that into s = 3a + 20 and try to work it out, and see how that goes.

OpenStudy (deercult):

That's also kind of a. Huge angle though. >n>;

OpenStudy (deercult):

So I think I def goofed somewhere along that line.

TheSmartOne (thesmartone):

mhmm, yeah you did you can plug in s = 3a + 20 in to s + a = 180 where we have s in the second equation, we can replace it with 3a + 20 since the first equation says that s is equal to 3a + 20 do you understand? :)

OpenStudy (deercult):

I'm gonna try that again w/ just the two variables. :V

OpenStudy (deercult):

OH. Yea, alright, geez, I was putting it in backwards. Silly me.

OpenStudy (joshtray):

Try solving for the s again

OpenStudy (deercult):

I'm gonna try to solve for a, actually! Gimme just one sec.

TheSmartOne (thesmartone):

yeah, if you substitute s = 3a + 20 in to the second equation, you'll be solving for a :)

OpenStudy (joshtray):

Use TheSmartOne's idea of substituting s with 3a+20

OpenStudy (joshtray):

That is if you want to solve for a

TheSmartOne (thesmartone):

at the end, we need to solve for both variables so it doesn't matter which one we start with :)

OpenStudy (deercult):

I got 26.65. That's a lot more reasonable than freakin', 330, at least.

OpenStudy (deercult):

Buuuuut it's wrong, still

OpenStudy (deercult):

B(

OpenStudy (joshtray):

What did you get for a?

OpenStudy (deercult):

Wait a tic, let me try something. I think I did some extra, unnecessary math, bc 26.65 is certainly Not it.

OpenStudy (deercult):

Nope, I was wrong either way ¯\_(ツ)_/¯

TheSmartOne (thesmartone):

can you show us your work? :)

OpenStudy (joshtray):

If the second equation is a+s=180, and you already know that s is 3a+20, put it as s in the equation a+s=180. Then work from there

OpenStudy (deercult):

Yea! (3a+20) + a = 180 -20 -20 (3a) + a = 180 /3a /3a a = 180/3 AND As I was typing that out I realized you can't... divide numbers like 180 by numbers w/ coefficients like that... (Right word? Who knows, not me.)

OpenStudy (joshtray):

Look you subtracted 20 from both sides right? 180-20 isn't 180

OpenStudy (deercult):

ER RIGHT. I meant 160.

OpenStudy (deercult):

I have that written down I swear. Sorry, I was trying to type really fast. Anyways, I ended up with 53.3... which I know isn't right bc that's not even, like. A Number. It repeats forever.

TheSmartOne (thesmartone):

and 3a + a = 4a so you need to divide both sides by 4 to isolate a :)

OpenStudy (joshtray):

You also should get to know that 3a+a=4a

OpenStudy (deercult):

OH That makes. A lot of sense

TheSmartOne (thesmartone):

3a + a is the same thing as a(3 + 1) = a(4) = 4a

OpenStudy (deercult):

OK, I got 40! When I plug it in like this: s= 3(40) + 20 s= 120 + 20 s= 140 And 140 + 40 is definitely 180. Success!

OpenStudy (deercult):

Also that's right. I sometimes forget that a number on its own is an implied 1. x')

TheSmartOne (thesmartone):

well done :D

OpenStudy (deercult):

Thank you both for all your help!

TheSmartOne (thesmartone):

It was my pleasure :)

OpenStudy (joshtray):

Your welcome